You have a battery marked \"6.00 V.\" When you draw a current of 0.225 A from it, the potential difference between its terminals is 5.09 V. What is the potential difference when you draw 0.487 A? THIS IS THE WHOLE QUESTION SOMEONE PLEASE HELP
Answer #1

Calula voltage is lost, for internal resistance of the battery

V_{lost}=V_{battery}- V_{measured}=6V-5.09V=1.18V

Applying Ohm's law to calculate the value of the resistance

R=V_{lost}/I=0.91V/0.225V=4.04\Omega

Ohm's law is again applied to calculate the difference of potential loss

For I=0.487A

V_{lost}=I\cdot R=0.487A\cdot 4.04\Omega =1.97V

the potenial difference between its terminals will be equal to

V_{measured}=V_{battery}-V_{lost} =6V-1.97V=4.03V